# The Rise of Ionized Gas Filaments in Early-Type Galaxies

### v1

@@ Abstract
Multiphase filamentary nebulae are ubiquitous in the brightest cluster galaxies (BCGs) of cool-core clusters, providing key insights into the cycle of baryons and the feeding and feedback of supermassive black holes. However, BCGs account for less than 1% of all early-type galaxies (ETGs). To broaden our understanding of how multiphase filamentary nebulae form in ETGs and connect to the greater picture of galaxy evolution, it is crucial to explore ETGs that are outside of the dense centers of galaxy clusters or groups. We present VLT-MUSE IFU observations of 126 nearby non-central ETGs, detecting warm ionized gas in 54 of them. 35 out of 54 display rotating disks that are morphologically and kinematically aligned with their stellar components, suggesting stellar mass loss as the origin of their warm gas. The remaining 19 host filamentary nebulae that are decoupled from the stellar components, resembling those observed in BCGs. These filamentary sources display unique emission line properties that cannot be fully explained by photoionization from post-asymptotic giant branch stars, active galactic nuclei, or fast gas shocks. For the eight filamentary sources that have been observed with Chandra, their soft X-ray emission indicates the presence of hot gas. We posit that their emission lines may be powered by EUV and X-ray radiation from the cooling of the hot gas, similar to cool-core clusters, but the detailed mechanisms and physical conditions may be different. As a detailed study case, we investigate NGC 4374, a non-central ETG with extensive Chandra observations, and find that its warm filaments are over-pressured compared to the hot filaments – opposite with what is observed in cool-core clusters.
@@ 1 Introduction
Early-type galaxies (ETGs), comprising elliptical and lenticular galaxies, were long believed to be devoid of gas species that are necessary for star-formation (SF). Their characteristically low star-formation rates (SFRs) led to the prevailing view that ETGs were passive, “red and dead” systems (Bregman, 1978; Mathews & Baker, 1971). However, such understanding of ETGs has been recently challenged due to the advent of telescopes with superior sensitivity and resolution. Gas that is multiphase in temperature and density, including (but not limited to) cold molecular ( K) (Edge, 2001; Salome & Combes, 2003; Young et al., 2011; Tremblay et al., 2018; Olivares et al., 2019), warm ionized ( K) (Heckman et al., 1989; Hatch et al., 2006; Hamer et al., 2016; Olivares et al., 2019), and diffuse hot ( K) gas (Sarazin, 1986; Fabbiano, 1989; Su et al., 2015) have been detected. The exciting detections of multiphase gas well motivate the astronomical community to revisit models describing the mass-assembly history of ETGs and their morphological evolution, given their minimal SFR.
@@
Observations of massive galaxies sitting in the center of galaxy clusters with sharply peaked X-ray surface brightness profiles (hereafter cool-core clusters) provide valuable insight into this puzzle. Their diffuse, hot intracluster medium (ICM) loses energy primarily through thermal bremsstrahlung as a result of free charges interacting with ions. For a spherical hot halo in quasi-hydrostatic equilibrium, there is a radius at which the cooling time (that is, the timescale for gas to radiate away its thermal energy content) is roughly the age of the system. Within this radius, radiative cooling drives an increase in gas density to maintain pressure support of the above gas layers. The increase in density is only plausible if there is an inward flow of gas funneling towards the center of the gravitational potential well, where typically the brightest cluster galaxy (BCG) resides. However, expected signatures such as X-ray coolant lines and high SFR are not observed. Thus, we are left with the “cooling-flow” problem which lies at the forefront of unresolved astrophysical puzzles (Lea et al., 1973; Fabian & Nulsen, 1977; Mathews & Bregman, 1978; Nulsen, 1986; Fabian, 1994).
@@
A heating source has been proposed to explain the lack of SF in these systems. An enticing candidate is the BCG’s active galactic nucleus (AGN), a highly energetic region driven by gas accretion onto the supermassive black hole (SMBH), which have typical masses of . In this scenario, some of the accreting rest-mass energy is released in radiation and kinetic outflows, and is enough to counter SF. Various lines of evidence suggest that cool-core clusters exhibit this phenomenon. For example, X-ray observations of the hot gas around BCGs in clusters with low core entropy ( 30 keV cm2) and short central cooling times ( 1 Gyr) reveal areas of significantly lower surface brightness (McNamara et al., 2000; Fabian et al., 2000; Wise et al., 2007; Cavagnolo et al., 2008; Dunn & Fabian, 2008). Radio observations find that these areas are also often filled with relativistic radio plasma (McNamara et al., 2005; Vantyghem et al., 2014). It is expected that AGN jets excavate these pockets of diffuse hot gas, and impart energy on their surroundings, thus preventing cooling and inhibiting SF. However, it is unclear how this mechanical energy is transferred to thermal energy. Aside from these clues provided by X-ray and radio observations, sub-mm and optical spectra show extended nebular emission in cool-core clusters (Edge, 2001; Salome & Combes, 2003; Cavagnolo et al., 2008; Rafferty et al., 2008; Olivares et al., 2019). Such findings have emphasized the role that multiphase emission line nebulae may play in explaining the connection between AGN feedback and SF.
@@
Atacama Large Millimeter Array (ALMA) and VLT Multi Unit Spectroscopic Explorer (MUSE) observations have revolutionized our understanding of the multiphase gas properties in massive ETGs. ALMA and MUSE provide a thorough view of the cold molecular gas (often traced by CO rotational lines) and warm ionized gas (typically traced by the H recombination line), respectively. The exceptional spatial and spectral resolution of these facilities allow their morphologies and kinematics to be viewed in great detail. In fact, ALMA observations of cool-core clusters find cold molecular gas in the form of filamentary networks, that are co-spatial and co-moving with the warm ionized gas filaments detected by MUSE (Tremblay et al., 2018; Olivares et al., 2019, 2022). This coupling between gas phases strongly suggest that they are linked. Condensation models suggest that the warm filaments envelope the cold molecular filaments, as these phases ultimately condense out of the thermally unstable hot halo. These multiphase nebulae often wrap around the aforementioned X-ray cavities, which could be explained by the bubbles buoyantly rising, and in doing so adiabatically lifting the low-entropy gas to radii where thermal instabilities are induced (Revaz et al., 2008; Li & Bryan, 2014).
@@
Studies of the most massive ETGs—such as brightest cluster galaxies (BCGs) and brightest group galaxies (BGGs; e.g., Olivares et al. (2022))—provide important but narrow insights into the formation of multiphase gas. BCGs represent only a small subset of the ETG population and occupy unique, extreme environments. They typically reside in the hot, dense cores of their host clusters’ intracluster medium (ICM). In these locations, the high ambient gas density makes it difficult to isolate and probe the truly low-entropy gas associated with cooling. Moreover, BCGs likely have complex evolutionary histories, often shaped by multiple mergers with surrounding satellites, further complicating the interpretation of observations. Extending spatially resolved studies to systems located away from the extreme centers of clusters and groups (hereafter “non-central” ETGs) will broaden our understanding of multiphase gas formation in ETGs and contribute to the larger picture of galaxy evolution.
@@
Eskenasy et al. (2024) analyzed the multiphase gas properties of 16 nearby, non-central ETGs selected from the ATLAS3D survey (Cappellari et al., 2011) using archival MUSE and Chandra X-ray observations. In this sample, the majority of sources with H emission have warm gas in the form of rotating disks, unlike the massive BCGs. Many of these rotating gas disks are kinematically coupled with the stars, suggesting that a significant amount of their warm gas may have formed from the shedding of their older stellar population. Only three systems in that sample show filamentary nebulae, and each of these three do not show coupling between their stars and gas. Interestingly, the Chandra observations reveal that the only sources with signatures of hot gas reservoirs were those with filamentary warm gas. They conclude that the observed H filaments likely condensed out of the thermally unstable hot phase, similar to BCGs, where warm gas filaments are believed to have cooled from their hot ICM. Although this study, as well as others (Temi et al., 2022), have paved the way for assessing multiphase gas formation and AGN feedback in non-central ETGs, the sample size is limited. Only two of the three ETGs with H filaments resemble the extended filamentary networks observed in BCGs. The literature lacks an adequate sample size of filamentary emission-line nebulae, to properly test the feasibility of cooling in typical ETGs.
@@ 2 Sample Description
In this paper we study 126 nearby ( Mpc) ETGs, a subset of the 50 Mpc Galaxy Catalog (50MGC; Ohlson et al. (2023)). 50MGC consists of 15,424 galaxies within 50 Mpc and is the first uniform and homogeneous source of general galaxy properties at this distance. The catalog’s self-consistent measurements (including morphological classification, mass, distance, etc.) allow for a broad range of analyses of local galaxies. The close proximity to these galaxies permits detailed observations of their stellar and gaseous components.
@@
In addition to these cuts made using the catalog’s published data, we manually removed 32 ETGs that were classified as either BGGs or BCGs in the literature. A large portion of removed BGGs are those within the Complete Local-volume Group Sample (O’Sullivan et al., 2017). Performing these cuts results in our sub sample comprised of 126 ETGs sitting outside of dense group and cluster cores, as well as some in truly isolated environments. This statistically significant sample of nearby, non-central ETGs is ideal for testing the internal physics driving multiphase gas formation across a broad stellar mass range (6.20 ).
@@ 3.1 MUSE Optical Integral Field Spectroscopy
Archival MUSE integral field unit (IFU) observations were used to deduce the nebular and stellar properties of our sample of ETGs. Table. LABEL:tab:muse provides the observation log for all sources in our sample. MUSE is a wide field of view (11′), image-slicing IFU spectrograph capable of achieving exquisite spatial resolution. The data were reduced using the European Southern Observatory Recipe Execution Tool (ESOREX v.3.13.1; Weilbacher et al. (2020)). The final datacube was sky subtracted using version 2.1 of the Zurich Atmosphere Purge (ZAP; Soto et al. (2016)) in addition to the data reduction pipeline’s internal sky subtraction method. The data are corrected for galactic extinction using the dust maps of Schlegel et al. (1998). The final MUSE datacube covers wavelengths between 4750-9300 Å.
@@
To remove noise from the emission line maps, we exclude pixels with a signal-to-noise ratio and gas velocity dispersions km s-1. Further pixels are removed if their 1 velocity dispersion errors exceed a given threshold. This cut-off was determined on a source-by-source basis by calculating the average velocity dispersion error in an offset region.
@@ 3.1.1 Stellar kinematics
Using the Galaxy IFU Spectroscopy Tool pipeline (GIST; Bittner et al. (2019)), the stellar kinematics were extracted from the MUSE datacube. The Voronoi tesselation routine outlined in Cappellari & Copin (2003) was used to spatially bin the IFU data. Spectra
were then logarithmically re-binned, followed by additional spatial binning. The pPXF routine (Cappellari & Emsellem, 2004) produced line-of-sight velocity and velocity dispersion maps for an initial velocity dispersion guess and rest-frame wavelength range. This wavelength range was determined on a source-by-source basis by searching for a region in the spectrum void of emission or absorption lines. Often times, the resultant stellar velocity and velocity dispersion maps contained foreground stars which were manually removed using a spatial mask.
@@ 3.2.1 X-ray luminosities
To extract the 0.5-2.0 keV luminosities for each source, we extracted counts within a circular region centered on the source with a radius of one effective radius (Re). We derived Re using the stellar mass-size relation given in Eq. 1 in Fernández Lorenzo et al. (2013). To convert our net count rates to energy fluxes, we used the PIMMS Mission Count Rate Simulator111https://cxc.harvard.edu/toolkit/pimms.jsp. We assumed a Plasma/APEC spectral model with solar metallicity and keV222We assume keV and keV for NGC 1266 and NGC 4374, respectively (Su et al., 2015). These priors are typical for the hot ISM of ETGs (Su & Irwin, 2013). We refer to Eskenasy et al. (2024) for describing how varying the metallicity and temperature impacts the X-ray luminosity using this spectral model. HI column densities were gathered from the HEASARC tool (Bekhti et al., 2016) to account for absorption along the line of sight.

### v2

@@ Abstract
Multiphase filamentary nebulae are ubiquitous in the brightest cluster galaxies (BCGs) of cool-core clusters, providing insight into baryon cycling and the feeding and feedback of supermassive black holes. However, BCGs account for less than 1% of all early-type galaxies (ETGs). To broaden our understanding of how multiphase filamentary nebulae form in ETGs and connect to the greater picture of galaxy evolution, it is crucial to explore ETGs that are outside of the dense centers of galaxy clusters or groups. We present VLT-MUSE IFU observations of 126 nearby non-central ETGs, detecting warm ionized gas in 62 of them. 35/62 host rotating gas disks with the majority of them morphologically and kinematically aligned with their stellar components, suggesting stellar mass loss may dominate their warm-gas origin. The remaining 27 host filamentary nebulae, often decoupled from the stellar components, resembling those observed in BCGs. These filamentary sources display unique emission line properties that cannot be fully explained by photoionization from post-asymptotic giant branch stars, active galactic nuclei, or fast gas shocks alone. For the twelve filamentary sources that have Chandra data, their soft X-ray emission exceeds or is consistent with (within uncertainties) unresolved low-mass X-ray binary emission, indicating that filamentary systems generally host an appreciable hot gas reservoir. We suggest that cooling-related processes, e.g., self-irradiation associated with the cooling hot gas, may contribute to the powering of warm gas line emission, similar to cool-core clusters, though the detailed mechanisms and physical conditions may differ. As a case study, we investigate NGC 4374, a non-central ETG with extensive Chandra observations, and find that its warm filaments are over-pressured compared to the hot filaments – opposite to what is observed in cool-core clusters.
@@ 1 Introduction
Early-type galaxies (ETGs), comprising elliptical and lenticular galaxies, were long believed to be devoid of gas species that are necessary for star-formation (SF). Their characteristically low star-formation rates (SFRs) led to the prevailing view that ETGs were passive, “red and dead” systems (Bregman, 1978; Mathews and Baker, 1971). However, such understanding of ETGs has been recently challenged due to the advent of telescopes with superior sensitivity and resolution. Gas that is multiphase in temperature and density, including (but not limited to) cold molecular ( K) (Edge, 2001; Salome and Combes, 2003; Young et al., 2011; Tremblay et al., 2018; Olivares et al., 2019), warm ionized ( K) (Heckman et al., 1989; Hatch et al., 2006; Hamer et al., 2016; Olivares et al., 2019), and diffuse hot ( K) gas (Sarazin, 1986; Fabbiano, 1989; Su et al., 2015) have been detected. The exciting detections of multiphase gas well motivate the astronomical community to revisit models describing the mass-assembly history of ETGs and their morphological evolution, given their minimal SFR.
@@
Observations of massive galaxies sitting in the center of galaxy clusters with sharply peaked X-ray surface brightness profiles (hereafter cool-core clusters) provide valuable insight into this puzzle. Their diffuse, hot intracluster medium (ICM) loses energy primarily through thermal bremsstrahlung as a result of free charges interacting with ions. For a spherical hot halo in quasi-hydrostatic equilibrium, there is a radius at which the cooling time (that is, the timescale for gas to radiate away its thermal energy content) is roughly the age of the system. Within this radius, radiative cooling drives an increase in gas density to maintain pressure support of the above gas layers. The increase in density is only plausible if there is an inward flow of gas funneling towards the center of the gravitational potential well, where typically the brightest cluster galaxy (BCG) resides. However, expected signatures such as X-ray coolant lines and high SFR are not observed. Thus, we are left with the “cooling-flow” problem which lies at the forefront of unresolved astrophysical puzzles (Lea et al., 1973; Fabian and Nulsen, 1977; Mathews and Bregman, 1978; Nulsen, 1986; Fabian, 1994).
@@
A heating source has been proposed to explain the lack of SF in these systems. An enticing candidate is the BCG’s active galactic nucleus (AGN), a highly energetic region driven by gas accretion onto the supermassive black hole (SMBH), which has a typical mass of . In this scenario, some of the accreting rest-mass energy is released in radiation and kinetic outflows, and is enough to counter SF. Various lines of evidence suggest that cool-core clusters exhibit this phenomenon. For example, X-ray observations of the hot gas around BCGs in clusters with low core entropy ( 30 keV cm2) and short central cooling times ( 1 Gyr) reveal regions of depressed surface brightness (McNamara et al., 2000; Fabian et al., 2000; Wise et al., 2007; Cavagnolo et al., 2008; Dunn and Fabian, 2008). Radio observations find that these areas are also often filled with relativistic radio plasma (McNamara et al., 2005; Vantyghem et al., 2014). It is expected that AGN jets excavate these pockets of diffuse hot gas, and impart energy on their surroundings, thus preventing cooling and inhibiting SF (though, the exact mechanism that thermalizes the atmosphere is unknown). Aside from these clues provided by X-ray and radio observations, sub-mm and optical spectra show extended nebular emission in cool-core clusters (Edge, 2001; Salome and Combes, 2003; Cavagnolo et al., 2008; Rafferty et al., 2008; Olivares et al., 2019). Such findings have emphasized the role that multiphase emission line nebulae may play in explaining the connection between AGN feedback and SF.
@@
Atacama Large Millimeter Array (ALMA) and VLT Multi Unit Spectroscopic Explorer (MUSE) observations have revolutionized our understanding of the multiphase gas properties in massive ETGs. ALMA and MUSE provide a thorough view of the cold molecular gas (often traced by CO rotational lines) and warm ionized gas (typically traced by the H recombination line), respectively. The exceptional spatial and spectral resolution of these facilities allow their morphologies and kinematics to be viewed in great detail. In fact, ALMA observations of cool-core clusters find cold molecular gas in the form of filamentary networks, that are co-spatial and co-moving with the warm ionized gas filaments detected by MUSE (Tremblay et al., 2018; Olivares et al., 2019, 2022). This coupling between gas phases strongly suggests that they are linked. Condensation models suggest that the warm filaments envelope the cold molecular filaments, as these phases ultimately condense out of the thermally unstable hot halo. These multiphase nebulae often wrap around the aforementioned X-ray cavities, which could be explained by the bubbles buoyantly rising, and in doing so adiabatically lifting the low-entropy gas to radii where thermal instabilities are induced (Revaz et al., 2008; Li and Bryan, 2014).
@@
Studies of the most massive ETGs—such as brightest cluster galaxies (BCGs) and brightest group galaxies (BGGs; e.g., Olivares et al., 2022)—provide important but narrow insights into the formation of multiphase gas. BCGs represent only a small subset of the ETG population and occupy unique, extreme environments. They typically reside in the hot, dense cores of their host clusters’ intracluster medium (ICM). In these locations, the high ambient gas density makes it difficult to isolate and probe the truly low-entropy gas associated with cooling. Moreover, BCGs likely have complex evolutionary histories, often shaped by multiple mergers with surrounding satellites, further complicating the interpretation of observations. Extending spatially resolved studies, using integral-field unit (IFU) spectroscopic observations, to systems located away from the extreme centers of clusters and groups (hereafter “non-central” ETGs) will broaden our understanding of multiphase gas formation in ETGs and contribute to the larger picture of galaxy evolution.
Indeed large IFU surveys such as ATLAS3D (Cappellari et al., 2011), CALIFA (Sánchez et al., 2012; Walcher et al., 2014), SAMI (Croom et al., 2012; Bryant et al., 2015), and MaNGA (Bundy et al., 2014) have explored the spatially resolved properties of nearby galaxies, including non-central ETGs. These surveys have transformed our knowledge of key properties that were largely inaccessible with single-slit spectroscopy, including but not limited to metallicity gradients, baryonic and dark matter distributions, and region-dependent ionization mechanisms. Despite this significant progress, these large surveys are not optimized to probe the detailed multiphase gas structure. In particular, their spatial resolution and wavelength coverage are generally insufficient to fully resolve filamentary warm gas structures and to exploit a complete set of optical emission-line-diagnostics. For example, while the SAURON spectrograph, used in the ATLAS3D survey, has detected warm ionized gas in NGC 4374 (Sarzi et al., 2006), its narrow field of view and small wavelength coverage (which excludes rest-frame H) preclude a truly detailed mapping of the nebulae. Spatially and spectrally resolving such structures is a necessary step toward understanding the role of cooling in producing filamentary multiphase gas formation in non-central ETGs, and can be accomplished with a facility like MUSE.
@@
Eskenasy et al. (2024) analyzed the multiphase gas properties of 16 nearby, non-central ETGs selected from the ATLAS3D survey using archival MUSE and Chandra X-ray observations. In this sample, the majority of sources with H emission have warm gas in the form of rotating disks, unlike the massive BCGs. Many of these rotating gas disks are kinematically coupled with the stars, suggesting that a significant amount of their warm gas may have formed from the shedding of their older stellar population. Only three systems in that sample show filamentary nebulae, and each of these three do not show coupling between their stars and gas. Interestingly, the Chandra observations reveal that the only sources with signatures of hot gas reservoirs were those with filamentary warm gas. This is consistent with the prevailing model in which the observed H filaments condense out of a thermally unstable hot phase, analogous to BCGs, where warm gas filaments are generally interpreted as having cooled from the hot ICM. Although this study, as well as others (Temi et al., 2022), have paved the way for assessing multiphase gas formation and AGN feedback in non-central ETGs, the sample size is limited. Only two of the three ETGs in Eskenasy et al. (2024) with H filaments resemble the extended filamentary networks observed in BCGs. The literature lacks an adequate sample size of filamentary emission-line nebulae, to properly test the feasibility of cooling in typical ETGs.
@@ 2 Sample Description
In this paper we study 126 nearby ( Mpc) ETGs, a subset of the 50 Mpc Galaxy Catalog (50MGC; Ohlson et al., 2023). 50MGC consists of 15,424 galaxies within 50 Mpc and is the first uniform and homogeneous source of general galaxy properties at this distance. The catalog’s self-consistent measurements (including morphological classification, mass, distance, etc.) allow for a broad range of analyses of local galaxies. The close proximity to these galaxies permits detailed observations of their stellar and gaseous components. Furthermore, the spatial resolution of MUSE allows us to explore the nebulae within our sample on sub-kpc scales. Fig. 1 compares the spatial resolution and wavelength coverage of our sample with that of the ATLAS3D, CALIFA, SAMI, and MaNGA IFU surveys - which have transformed our understanding of the gaseous properties of galaxies. Our sample can build on this progress by combining enhanced spatial resolution with an average spectral resolution , sufficient for identifying key, individual diagnostic emission lines.
@@
In addition to these cuts made using the catalog’s published data, we manually removed 32 ETGs that were classified as either BGGs or BCGs in the literature. A large portion of removed BGGs are those within the Complete Local-volume Group Sample (O’Sullivan et al., 2017). Performing these cuts results in our sub sample comprised of 126 ETGs sitting outside of dense group and cluster cores, as well as some in truly isolated environments. This statistically significant sample of nearby, non-central ETGs is ideal for testing the internal physics driving multiphase gas formation across a broad stellar mass range (6.20 ). The top panel of Fig. 2 shows the stellar mass distribution of all ETGs as well as late-type galaxies (LTG) within the 50MGC sample (with available stellar mass estimates). The overlaid solid black histogram shows the stellar mass distribution of our subsample described above.
@@ 3.1 MUSE Optical Integral Field Spectroscopy
Archival MUSE IFU observations were used to deduce the nebular and stellar properties of our sample of ETGs. Table. LABEL:tab:muse provides the observation log for all sources in our sample. MUSE is a wide field of view (11′), image-slicing IFU spectrograph capable of achieving exquisite spatial resolution. The data were reduced using the European Southern Observatory Recipe Execution Tool (ESOREX v.3.13.1; Weilbacher et al., 2020). The final datacube was sky subtracted using version 2.1 of the Zurich Atmosphere Purge (ZAP; Soto et al., 2016) in addition to the data reduction pipeline’s internal sky subtraction method. The data are corrected for galactic extinction using the dust maps of Schlegel et al. (1998). The final MUSE datacube covers wavelengths between 4750-9300 Å.
@@
To remove noise from the emission line maps, we exclude pixels with a signal-to-noise ratio and gas velocity dispersions km s-1. Further pixels are removed if their 1 velocity and velocity dispersion uncertainties exceed a given threshold. This cut-off was determined on a source-by-source basis by calculating the average velocity dispersion uncertainty in an offset region.
@@ 3.1.1 Stellar kinematics
Using the Galaxy IFU Spectroscopy Tool pipeline (GIST; Bittner et al., 2019), the stellar kinematics were extracted from the MUSE datacube. Spaxels were tessellated with the Voronoi SNR method outlined in Cappellari and Copin (2003), to a target SNR40 per bin. For galaxies with low surface brightness, this target could not be met and thus was lowered accordingly down to SNR15. Spectra
were then logarithmically re-binned in wavelength. The pPXF routine (Cappellari and Emsellem, 2004) produced line-of-sight velocity and velocity dispersion maps for an initial velocity dispersion guess and rest-frame wavelength range. This wavelength range was determined on a source-by-source basis by searching for a region in the spectrum void of emission or absorption lines. Foreground stars were manually excluded using a spatial mask on the final maps.
@@ 3.2.1 X-ray luminosities
To extract the 0.5-2.0 keV luminosities for each source, we extracted counts within a circular region centered on the source with a radius of one effective radius (Re). We derived Re using the stellar mass-size relation given in Eq. 1 in Fernández Lorenzo et al. (2013). To convert our net count rates to energy fluxes, we used the PIMMS Mission Count Rate Simulator111https://cxc.harvard.edu/toolkit/pimms.jsp. We assumed an absorbed plasma APEC model, phabsapec, with solar metallicity and keV222We assume keV and keV for NGC 1266 and NGC 4374, respectively (Su et al., 2015). These priors are typical for the hot ISM of ETGs (Su and Irwin, 2013). We refer to Eskenasy et al. (2024) for describing how varying the metallicity and temperature impacts the X-ray luminosity using this spectral model. HI column densities were gathered from the HEASARC tool (Bekhti et al., 2016) to account for absorption along the line of sight.